News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal (opens in new window)
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 16 December 2016 9.00am EST

During a real-time pass on Dec 9 the ACIS Digital Processing Assembly Side-A (DPA-A) was detected in an unexpected powered-off state. Engineering assessment indicated behavior consistent with prior cases where the DPA powered-down. In each of those cases a fault-tree analysis concluded the event was due to a single event upset. The DPA was powered back on and the required software patches reloaded during a real-time pass on Dec 9. Recovery was complete at 7:13pm EST on Dec 9 with 29.1ks of science loss. The scheduled observation of Abell 1795 Outskirts was impacted by the DPA-A shut-down and the lost science time will be rescheduled in future weeks.

Of note this week was the release on Dec 15 of the Cycle 19 Call for Proposals. The proposal submission deadline is 15 March 2017. In conjunction with the call was the release of the CIAO 4.9 analysis software system, CALDB 4.7.3, MARX 5.3.2, DS9 7.5 and of CXCDS 10.5.3 for standard processing. For details see: http://cxc.harvard.edu/proposer/ and http://cxc.harvard.edu/ciao/releasenotes/ciao_4.9_release.html

The schedule of targets for the next week is shown below and includes an observation of HS 0810+2554 coordinated with NuSTAR.



------------------------------------------
RXJ1131-1231 ACIS-S Dec 19 CygnusA-WHotspot ACIS-I 3C129.1 ACIS-S Radiation Belts Dec 20 Abell1795outskirts ACIS-I IGRJ00370+6122 HRC-I Dec 21 3C129.1 ACIS-S 2MASSJ06170531+8353 ACIS-S IRAS17020+4544 HRC-S /LETG Q2237+0305 ACIS-S Dec 22 Radiation Belts Dec 23 HS0810+2554 ACIS-S NGC4419 ACIS-S Dec 24 IRAS17020+4544 HRC-S /LETG Abell1795outskirts ACIS-I Radiation Belts Dec 25

------------------------------------------

All spacecraft subsystems continued to support nominal operations.

Return to Status Reports